OpenStudy (anonymous):

solve the equation 6 |6-4x| = 8x+4

OpenStudy (yttrium):

It can be written as |36-24x| = 8x+4 and still, you need to split this into two cases 35-24x = 8x+4 and -(35-24x) = 8x+4 Just evaluate for x.

OpenStudy (anonymous):

11x?

OpenStudy (yttrium):

It is 35-4 = 24x+8x x = 31/32

OpenStudy (anonymous):

ow i see

OpenStudy (yttrium):

and -35+24x = 8x+4 whic is equal to 16x = 39 x = 39/16

OpenStudy (anonymous):

thank you so much:))

OpenStudy (anonymous):

a. x=1 b. x=5/2 or x=2 c. x=5/2 d. x=5/2 or x=1

OpenStudy (yttrium):

Oww. Let's do this again. |6-4x| = (8x+4)/6 split this into two cases(1st) 6(6-x) = 8x+4 36-24x = 8x+4 32 = 32x x = 1 and (2nd case) -6(6-4x) = 8x+4 -36+24x = 8x+4 16x = 40 x = 5/2 Then we get x= 1 and x = 5/2

OpenStudy (anonymous):

thanks again :)